# Histopathology Slide Indexing and Search: Are We There Yet?

### v1

@@ 4 Discussion
In this case study, we evaluated the performance and clinical utility of state-of-the-art histopathology slide search engines, Yotixxel, SISH, and RetCCL on real patient samples. While these models demonstrate significant advancements in the field, all exhibit certain shortcomings when trialing them on real patient cases deemed to be of moderate diagnostic difficulty by pathologists, indicating further progress is warranted.
We found that RetCCL was the most reliable in tissue retrieval by correctly identifying within the first five retrieved slides the correct tissue of diagnosis for all three cases. In contrast, both Yottixel and SISH failed to retrieve the correct tissue of origin at all top five ranked positions for two out of the three cases, which was unexpected given prior work demonstrating the high performance of these models.
For subtype search, we found overall higher performance amongst all three models than for tissue search, a discrepancy which has been reported previously. However, a significant component of this higher performance may be attributed to the narrowing of the search domain as all three subtype search methods require pre-selecting the tissue of interest. We also found that commonly used metrics in the literature for evaluation, such as mMW@5, may also be less clinically useful for real life scenarios. For example, while SISH subtype search on Slide 1 correctly retrieves a LUAD slide, when using the mMW@5 metric, this would still result in a perfect score of 100, despite the exclusion of the other 67 slides pertaining to LUAD in our dataset.
At the patch level, we find that RetCCL was more capable of identifying similar patches containing lung glandular cells but all three models struggled to differentiate between more subtle features of malignancy, such as the loss of cell borders and prominent nucleoli. These differences in performance may be attributed to RetCCL’s use of self-supervised learning for feature representation over the use of previous methods. Similarly, we also found that RetCCL outperformed both SISH and Yotixxel in retrieving patches containing normal alveolar tissue. Of note, the utility of these models in retrieving cancerous versus non-cancerous slides has yet to be formally investigated as these search engines were not intentionally designed for this task. Nevertheless, for search results to be accepted as reliably accurate by the clinical community, particularly on tasks such as patch-level retrieval where differences can be more pronounced, these capabilities may need to be eventually demonstrated as well.
As the field of digital pathology continues to evolve, we anticipate exciting developments in the near future. These will likely include more efficient and reliable systems for indexing and searching of histopathology slides, increasingly robust algorithms for feature extraction, and potentially transformative diagnostic tools. However, given the high-stakes nature of patient care, we anticipate a significant amount of work ahead to ensure the validity and generalizability of these models prior to clinical adoption.
To this end, we propose a framework of minimal requirements designed to facilitate the development of these systems in a manner that is both fair and transparent while minimizing patient harm and maximizing clinical utility:
@@
    Open access to indexed databases and code: Achieving clinical readiness for these systems demands the sharing of both model development code and indexed databases. This transparency not only fosters scientific integrity but also accelerates the pace of innovation through collective efforts.
@@
    Replicability and generalizability of results across different health systems: A vital measure of clinical applicability is the model’s performance across diverse clinical populations and environments, as in the case of real-world health systems.
@@
    Clear explanation of algorithmic innovation and clinical benefit: For clinical adoption, the unique aspects of each model and their direct value need to be clearly communicated. We propose the inclusion of two additional tasks where image search will be most helpful to pathologists in identifying the tissue of cancerous origin (in the metastatic setting) and distinguishing between cancerous and non-cancerous tissue.
@@
    Fair and Balanced Comparison of Models: Clinical adoption of these systems is contingent upon their measured performance relative to their peers. This requires fair and balanced comparisons using identical pre- and post-processing methods and the adoption of new metrics that accurately reflect clinical utility.
@@
    Realistic and Transparent Discussion of Computational Efficiency: Clinical applicability not only requires theoretical efficiency, but also hinges on the performance of these systems in a real-world, high-demand environment. Thus, discussions on computational efficiency should reflect the realities of clinical implementation.
As we continue to make strides in the development of histopathology slide search engines, it is imperative that we maintain a steadfast focus on these requirements. By doing so, we will ensure that these systems are not only theoretically sound, but also ready for meaningful clinical application.
@@
The test slides along with the updated source codes for all three methods used to generate the results can be found at github.com/jacobluber/PathologySearchComparison. All the data used in databases are publicly available at portal.gdc.cancer.gov. The list of data included in the database can also be found in the github repository above.
@@
Supplementary Methods
@@
Search Engines Used in Comparison
@@
RetCCL, drawing inspiration from both Yottixel and SISH, adopts a distinct approach. Instead of clustering patches based on RGB histogram values, RetCCL first obtains contrastive-based feature vectors for each patch within the segmented foreground tiles. These features serve as inputs for a 9-class K-means clustering. Within each cluster, an additional K-means process based on spatial coordinates is performed to select 20% of the patches. These selected patches form the mosaics, which constitute the database. The proposed feature extraction algorithm in RetCCL utilizes a clustering-guided contrastive learning method, employing the InfoNCE loss introduced in [9] (Fig. 1c). Given the prevalence of normal cells in WSIs, learning irregularities from a limited number of patches becomes crucial. The self-supervised feature extractor employs two InfoNCE losses to capture irregular regions in patches.
@@
Supplementary Results
Lung Adenocarcinoma Slide Retrieval
Slide 1 is a Lung Adenocarcinoma (LUAD) case from a patient’s partial lobectomy. The highest performing model in the literature on retrieving LUAD slides using subtype search is SISH; in their seminal paper, Chen et. al reported a mMV@5 (the mean average precision for the top 5 results) of 70-80 on retrieving LUAD from TCGA when applying subtype retrieval, which was maintained when testing against an independent test cohort of 1,377 LUAD slides. In contrast, the authors of RetCCL reported a mMV@5 of 84.01 relative to Yottixel’s 70.96 on LUAD when applying subtype retrieval on 538 FFPE WSIs.
Amongst the three models tested, only RetCCL WSI organ search was able to retrieve slides that were of lung origin (Fig. 2a). Yottixel organ search only identifies slides pertaining to brain tissue, despite the presence of over 100 lung tissue slides in our dataset. We also found poor performance when using SISH organ retrieval as only one slide was retrieved corresponding to liver tissue. For subtype retrieval, we found that all three methods were now able to correctly retrieve several slides of LUAD in the top ranked position by subtype search. The extended results are provided in Extended Results.
Low Grade Glioma Slide Retrieval
We next investigated the performance of these algorithms on a Low Grade Glioma (LGG) case in Slide 2. This case was chosen given the known difficulty of differentiating Lower from Higher Grade Gliomas by morphology alone with molecular-based studies recently replacing histological classification as the gold standard for subtyping and grading.
The highest performing algorithm in the literature for LGG retrieval by subtype search is again SISH, which achieved a mMV@5 of nearly 100 when using the TCGA dataset’s 838 LGG slides for subtype search, which is significantly higher from Yottixel’s mMV@5 of 80-90. In contrast, the authors of RetCCL reported lower performance relative to Yottixel on LGG retrieval by subtype search on an independent cohort of 1699 brain FFPE WSIs, achieving a mMV@5 of 83.73 versus 89.77, respectively.
For Slide 2, we found that only RetCCL correctly retrieves LGG slides on subtype search whereas both Yottixel and SISH retrieve Glioblastoma (GBM) cases at all top five positions (Fig. 2b). For organ retrieval, Yottixel, SISH, and RetCCL all correctly retrieve slides containing brain tissue at top ranking positions with comparable performance. The extended results can be found in Extended Results.
Hepatocellular Carcinoma Slide Retrieval
Slide 3 is from a patient with Hepatocellular Carcinoma (LIHC) who underwent a liver biopsy by fine needle aspiration. Chen et. al previously demonstrated comparable performance on subtype retrieval for LIHC between SISH and Yottixel with a mMV@5 of 80-90 on the TCGA database. The authors of RetCCL also achieved similar performance as Yottixel across an independent cohort of 398 LIHC WSIs with a mMV@5 of 96.06 and 93.65, respectively. In regards to liver organ search, RetCCL has shown higher performance, achieving a mMV@10 of 89.97 versus 63.75 for Yottixel on 628 FFPE liver WSIs. In contrast, SISH and Yottixel achieved similar performance on liver tissue retrieval with a mMV@10 of 60-70 using TCGA.
As with lung tissue, we found poor performance for both Yotixxel and SISH on liver tissue retrieval as both queries resulted in brain slides only (Fig. 2c). We again find higher performance with RetCCL on organ search, which achieves perfect performance in identifying liver slides at all top five ranked positions. In contrast, for subtype search, we find higher performance for Yottixel and SISH, which now correctly only retrieve LIHC cases, whereas for RETTCL, several cases of Cholangiocarcinoma (CHOL) are mistakenly retrieved. The extended results are provided in Extended Results.
Patch Retrieval on Lung Adenocarcinoma
To better interpret granular differences in feature extraction, we next evaluated the performance of the three models on identifying similar patches of interest. For this task, we chose a patch from Slide 1 corresponding to an area of neoplastic LUAD cells with expected features of malignancy, such as abundant cytoplasm, nuclear crowding, loss of clear intracellular borders, and prominent nuceloi.
At the patch level, we find RetCCL is the most consistent at retrieving patches of lung glandular cell. However, these patches only contain normal glandular cells whereas in our queried patch, we have an abundance of neoplastic glandular cells, suggesting a weakness in discriminating more granular, cellular features of malignancy, which are subtle yet still discernible to pathologists. For Yottixel and SISH, we see less consistent retrieval of lung glandular tissue with several examples of primarily stroma and scant glandular cells as well as several examples of normal glandular tissue. As a whole, this suggests that RetCCL may have better performance in capturing higher-level features but all three models may have difficulties in discriminating more granular, subtle features of malignancy.
We also queried a patch containing normal alveolar tissue, which has more visually pronounced structural differences from glandular tissue, including a network of thin epithelial cell walls. We now find that all three models now correctly retrieve patches corresponding to normal alveolar tissue, although some patches from SISH and Yottixel include thickened layers of fibroblasts interspaced within the alveoli walls, which can be reactive in the setting of inflammation or malignancy. This is in contrast to queried slide where the alveoli walls only contain a single layer of flattened squamous epithelial cells, which can also be seen in the patches retrieved by RetCCL.
Unprocessed Slides
@@
[ht] Unprocessed slides in the database for each model. Model Slide IDs Yottixel 9ecf91d4-0d9e-4400-bf38-99420acd14cc f18b6fc0-6f40-4f0d-82ef-0b092a21b6bf 846087b8-f70c-4970-a1b7-24d403229801 c95681f3-53d4-4b15-833d-ff68f171965e 71dc7ba0-a623-4aaf-9502-f2fe9d188401 dc5d0b4-04ff-4731-bda1-8ad7cd0fa345 SISH 2dc5d0b4-04ff-4731-bda1-8ad7cd0fa345 RetCCL All slides in the database processed.

### v2

@@ 4 Discussion
In this case study, we evaluated the performance and clinical utility of state-of-the-art histopathology slide search engines, Yotixxel, SISH, and RetCCL and HSHR. To our knowledge, this is the first independent, external validation of these four models. While all models demonstrate significant advancements in the field, we also noticed clinically significant shortcomings.
Based on our findings, we propose a framework of minimal requirements to facilitate the development of these systems both fairly and transparently while minimizing patient harm and maximizing clinical utility:
@@
    Richness of feature representations: We observed several clinically relevant inconsistencies across our queried and retrieved patches, suggesting the need for improvements in feature extraction. Our ablation study also highlights the importance of encoder architecture on model performance. Large pre-trained models such as Virchow ([23]), with extensive training on 1.5 million images, are a potential solution, similar to natural image processing models such as VGG16 ([19]) and ResNet-50 ([7]). This strategy can also facilitate the prioritization of downstream tasks while alleviating extensive computational burdens in the pre-training stage.
@@
    Systematic and rigorous evaluations: We found it challenging to evaluate relative model performance due to variable results across the four algorithms in our experiments. For example, RetCCL had the highest performance on brain tissue retrieval but Yottixel had better performance on breast tissue. Likewise, while we found that pathologists tended to rank patches retrieved by RetCCL highly, due to large inter-observer variation on quality, relative performance was still uncertain. Our findings highlight the need for systematic and rigorous methodologies for model evaluation.
@@
    Robustness: A vital measure of clinical applicability is model performance across diverse clinical populations and environments, as in the case of real-world health systems. This is motivated by our findings that all models experienced a notable decrease in performance when tested against unseen datasets. However, we were pleased to find that batch level effects from external factors such as microscope type did not result in statistically significant differences on tissue and subtyping performance. We propose that all future models be validated for their generalizability in addition to precision.
@@
    Replicability: We found variable degrees of replicability across the models evaluated in this work. Some were constructed for specific slide ratios and resolutions, while others were capable of handling variable slide formats. Models also differed in terms of ease of applicability and transparency due to differences in the availability of model development code and indexed databases. Model transparency fosters scientific integrity and accelerates the pace of innovation through collective efforts and unbiased evaluations.
@@
    Clinical benefit: For clinical adoption, the unique aspects of each model and their value need to be clear for end-users. To date, the ongoing development of WSI search algorithms has focused on achieving stepwise improvements in performance on tissue and subtyping search. While these are worthy goals, we propose the inclusion of new tasks based on existing challenges in the field of pathology, such as helping to determine the tissue of cancerous origin in the metastatic setting.
  6. 6.
    Computational Efficiency: Clinical applicability not only requires theoretical efficiency, but also hinges on the performance of these systems in a real-world, high-demand environment. Thus, discussions on computational efficiency should reflect the realities of clinical implementation. Especially, we need to make sure the processes of querying an expanding the database are really efficient. These two processes are of utmost importance for a sustainable search engine. We recommend the database indexing algorithms to be online (i.e. adding a new data point does not require compiling the whole database from scratch). Also, one shortcoming of models like RetCCL, SISH, and Yottixel is that they generate mosaics as a percentage of number of patches in a cluster. Now, if like GBM UPENN, slides become very large, number of patches becomes unnecessarily large and query time becomes very long. We recommend models such as HSHR at use a fixed size for patches for this matter.
As the field of digital pathology continues to evolve, we anticipate exciting developments in the near future. These will likely include more efficient and reliable systems for indexing and searching of histopathology slides, increasingly robust algorithms for feature extraction, and potentially transformative diagnostic tools. Given the high-stakes nature of patient care, we anticipate a significant amount of work ahead to ensure the validity of these models prior to clinical adoption. As we continue to make strides in the development of histopathology slide search engines, our proposed criteria ensures that these systems are not only theoretically sound but also ready for meaningful clinical adoption.
@@
The test slides along with the updated source codes for all three methods used to generate the results can be found at github.com/jacobluber/PathologySearchComparison. All other data used in databases are publicly available at portal.gdc.cancer.gov and www.cancerimagingarchive.net/. The list of data included in the database can also be found in the github repository above.
@@
Acknowledgements
We thank the following for their contributions: David Fernandez-Hazoury, M.D., Department of Pathology, Harbor-UCLA Medical Center; Kenechukwu Ojukwu, M.D., M.P.P.,Bone/Soft Tissue Fellow, Department of Pathology and Laboratory Medicine, National Clinician Scholars Program (NCSP) Fellow; Alex Oliveira-Kowaleski, M.D., UCLA Department of Pathology and Laboratory Medicine; Julie Y. Kim, D.O., M.S., Harbor-UCLA Medical Center; Daniel P. Stefanko, M.D. Ph.D., UCLA Department of Pathology and Laboratory Medicine.
@@ Acknowledgements
Datasets
{stable}
[!h] Summary of database used for comparison experiments. Abbreviations are based on [11]. Primary Site Project Name (Subtype) Abbr. Num. Slides Num. Selected Slides Brain Glioblastoma Multiforme GBM 2040 61 Brain Lower Grade Glioma LGG 1543 69 Lymphoid Neasm Diffuse Large B-cell Lymphoma DLBC 4 0 Breast Breast Invasive Carcinoma BRCA 2704 72 Lymphoid Neasm Diffuse Large B-cell Lymphoma DLBC 2 0 Bronchus and lung Lung Adenocarcinoma LUAD 1359 68 Lung Squamous Cell Carcinoma LUSC 1265 57 Mesothelioma MESO 2 0 Colon Colon Adenocarcinoma/Rectum Adenocarcinomaa COAD/READ 1307+18 59 (COAD) Lymphoid Neasm Diffuse Large B-cell Lymphoma DLBC 6 0 Sarcoma SARC 4 0 Liver and intrahepatic bile ducts Liver Hepatocellular Carcinoma LIHC 778 72 Cholangiocarcinoma CHOL 80 50
  • a
    Although from pathologist point of view Colon Adenocarcinoma and Rectum Adenocarcinoma are genetically and morphologically the same entity, TCGA considers them different projects.
{stable} {stable}
[!h] Unprocessed slides in the database for each model. Model Slide IDs Yottixel 9ecf91d4-0d9e-4400-bf38-99420acd14cc f18b6fc0-6f40-4f0d-82ef-0b092a21b6bf 846087b8-f70c-4970-a1b7-24d403229801 c95681f3-53d4-4b15-833d-ff68f171965e 71dc7ba0-a623-4aaf-9502-f2fe9d188401 2dc5d0b4-04ff-4731-bda1-8ad7cd0fa345 SISH 2dc5d0b4-04ff-4731-bda1-8ad7cd0fa345 RetCCL All slides in the database processed. HSHR f2d5aa37-d9ce-4264-a447-fc69dd0d7d85 a2658e39-e476-44b2-99ee-118056cf6201 f84130fe-4853-4252-a292-9372aeea4a5d 22904f9d-0788-463c-9961-02629cf9a85f
@@
Search Engines Methods
@@
RetCCL, drawing inspiration from both Yottixel and SISH, adopts a distinct approach. Instead of clustering patches based on RGB histogram values, RetCCL first obtains contrastive-based feature vectors for each patch within the segmented foreground tiles. These features serve as inputs for a 9-class K-means clustering. Within each cluster, an additional K-means process based on spatial coordinates is performed to select 20% of the patches. These selected patches form the mosaics, which constitute the database. The proposed feature extraction algorithm in RetCCL utilizes a clustering-guided contrastive learning method, employing the InfoNCE loss introduced in [17] (Fig. 1c). Given the prevalence of normal cells in WSIs, learning irregularities from a limited number of patches becomes crucial. The self-supervised feature extractor employs two InfoNCE losses to capture irregular regions in patches. It worth mentioning that RetCCL was not able to perform the indexing on UPENN GBM dataset within a reasonable run time and was not utilized for certain experiments involving this dataset.
@@
In HSHR, the first step is to train the ResNet18 [8] backbone encoder using the SimCLR [2] approach. Unlike the other methods, authors had not provided their backbone pre-trained weights, so we trained it from the scratch. The training data for this backbone was approximately $508\times 100=50,800$ randomly patches of size $224\times 224$. We trained it for 200 epochs using the same hyperparameters as the authors on 2 Nvidia A100 GPUs. Once the backbone is trained, it is used to extract the features for all densely-patched patches for each WSI. These features are used to train a 20-class k-means clustering algorithm. The features of the centroids of these clusters create the mosaic for each WSI. These features are then passed to CaEncoder for generating the hashes and attention weights. Unlike the backbone, the authors had provided the weights for their CaEncoder, and we used the same weights in our experiments. The outputs of CaEncoder is further used to create a hypergraph for the database using Eq. 15 in [14]. We used $K=10$ in this equation.
Once the hypergraph for database is constructed, each query slide would go through the same pipeline and becomes and turns into hash codes and attention weights. Using these values, the query can be appended to the hypergraph as a new vertex and a new hyperedge. Updating the hypergraph, the similarity score between this vertex and all vertices in the database can be calculate using Eq. 19 in [14]. Then the top-k results are returned. We used $\alpha=\beta=1$ in this equation.
Methodological Breakdown of Query and Retrieval Processes
Input: Image $I$
Output: Top 5 retrieved slides
1 Patch the image densely to get patches $p_{1},p_{2},\ldots$;
2 Perform RGB histogram clustering;
3 Perform spatial clustering and calculate mosaic patches;
4 Feed mosaic patches to KimiaNet for feature extraction and calculate barcode for each patch;
5 foreach patch in input slide do
6       Calculate hamming distance between barcode of input patch with barcodes of patches from all slides in database;
7
8Choose the median of the list of minimum hamming distances for each slide in database;
9 Retrieve the slides with the top five smallest medians;
Algorithm 1 Yottixel Algorithm
Input: Image $I$
Output: Similar slides to query slide $I$
1 Patch the image densely to get patches $p_{1},p_{2},\ldots$;
2 Perform RGB histogram clustering;
3 Perform spatial clustering and create mosaic patches;
4 Feed mosaic patches to DenseNet and VQ-VAE encoder to calculate parameters $h$ and $m$ respectively, and create the VEB tree;
5 Apply guided-search algorithm to tuples of $m$ and $h$ to calculate corresponding set of tuples $r$;
6 Create a set of candidate indices $mi,c+$ and $mi,c-$ along with the original $mi$;
7 Call helper functions forward-search and backward-search on $mi,c+$ and $mi,c-$ respectively;
8 Take the results $RI=\{r_{1},r_{2},\ldots,r_{k}\}$ from Guided-Search as input by Results Ranking Algorithm;
9 Return similar slides to query slide $I$;
Algorithm 2 SISH Algorithm
Input: Image $I$
Output: Top $k$ similar WSIs
1 Patch the image densely to get patches $p_{1},p_{2},\ldots$;
2 Feed patches to feature extraction algorithm;
3 Cluster based on extracted features, then on coordinates to create mosaic patches;
4 foreach patch in query WSI do
5       Perform Knn search to retrieve a bag of most similar patches in database to each patch, using cosine similarity in pretrained SSL encoder’s learned embedding space;
6
7Calculate entropy within each bag, reorder bags by entropy;
8 Remove bags with low quality based on mean of cosine similarity scores in top-5;
9 foreach bag do
10       Perform voting for each diagnosis within the bag, get the top-5 samples, then do majority vote to get associated WSI;
11
12Retrieve top-$k$ similar WSIs;
Algorithm 3 RetCCL Algorithm
Input: Image $I$
Output: Top $k$ similar WSIs
1 Patch the image densely to get patches $p_{1},p_{2},\ldots$;
2 Feed patches to feature extraction algorithm;
3 Cluster based on extracted features, then on coordinates to create mosaic patches;
4 foreach patch in query WSI do
5       Create a bag containing the query patch and its retrieved patches;
6
7Calculate entropy within each bag, reorder bags by entropy;
8 Remove bags with low quality based on mean of cosine similarity scores in top-5;
9 foreach bag do
10       Perform voting for each diagnosis within the bag, get the top-5 samples, then do majority vote to get associated WSI;
11
12Retrieve top-$k$ similar WSIs;
Algorithm 4 HSHR Algorithm
Complexity Analysis
The time complexities of the key operations in different algorithms are summarized in Table Complexity Analysis. In the table, $n$ (Yottixel) denotes the length of the hamming vector, $m$ (Yottixel) denotes the maximum number of patches for each WSI, $T$ (common across Yottixel, HSHR, and RetCCL) denotes the total number of slides in the database, $B$ (SISH) denotes the number of patches in a WSI, $K$ (RetCCL, HSHR) denotes the number of patches of the query slide, $M$ (RetCCL, HSHR) denotes the total number of diagnoses in the database, and $H$ (HSHR) denotes the hash vector size.
@@ 8 Remove bags with low quality based on mean of cosine similarity scores in top-5;
[ht] Time complexity analysis for different components of the compared methods. Algorithm Operation and Time Complexity Yottixel Hamming distance calculation: $O(n.T.m^{2})$ Minimum: $O(T.m^{2})$ Median: $O(T.\log(T))$ SISH Search Performance: $O(1)$ Ranking: $O(B’)$ ’ ), $B’=0.05\cdot B$ RetCCL Cosine similarity calculation: $O(K.B)$ Probability calculation: $O(M.K.B)$ Entropy calculation: $O(K.M)$ Sorting bags based on entropies: $O(K^{2})$ Mean of cosine similarity scores: $O(T.K)$ Removing bags with low quality: $O(K)$ HSHR Hamming distance: $O(H.T.M^{2})$ Sorting: $O(M)$ Incidence Matrix calculation: $O(K.M.T)$ Cross similarity: $O(T^{2})$ Vertex similarity: $O(T^{3})$ Hyperedge similarity: $O(T^{3})$ Final sorting: $O(T)$
Studies
Performance Metrics
As discussed, we use majority voting and average precision at $k$ as main performance metrics in different experiments. Supplementary Algorithms algorithm 5 and algorithm 6 summarize they way we defined these metrics. The important point about majority voting at $k$ is that if it returns None, that sample would not be counted in the average, while 0 outputs are counted towards average.
Data: $row$, $k$
Result: Result of the majority vote
1 $votes\leftarrow\text{empty list}$ for $i=1$ to $k$ do
[…]
3             append $-1$ to $votes$
4      else
5             append $ret\_site$ to $votes$
6
7if not $votes$ then
8       return 0
[…]
10       return 1
11else
12       if $most\_common[0][0]=-1$ then
13             return None
14      else
15             return 0
16
Algorithm 5 Majority Voting at k
Data: $row$, $k$
Result: Average precision at k
1 $relevant\_count\leftarrow 0$ $precision\_sum\leftarrow 0$ for $i=1$ to $k$ do
[…]
3             continue
[…]
5             $relevant\_count\leftarrow relevant\_count+1$ $precision\_sum\leftarrow precision\_sum+\frac{relevant\_count}{i}$
6
7if $relevant\_count=0$ then
8       return 0
return $\frac{precision\_sum}{\text{min}(k,relevant\_count)}$
Algorithm 6 Average Precision at k
Reader study
Seven pathologists were shown the top three ranked results on patch-level retrieval for one patch across three different H&E slides. All pathologists were shown the original queried patch of interest but were blinded to the algorithms involved, ranked order of the retrieved patches, and diagnoses of both the queried and retrieved patched. Pathologists were then asked for their Mean Opinion Score (MOS) based on their perspective on the quality of the results, ranked from one to five with higher scores indicating higher quality.
{sfigure}
[!ht] [Uncaptioned image] Reader Study: Average quality ratings and standard deviation for each algorithm as evaluated by seven pathologists.
{stable}
[!ht] Evaluation results of different methods on Reader Study slides for primary site retrieval task. Method Reader Study MV@1 MV@3 MV@5 MV@10 AP@3 AP@5 YOTTIXEL + KimiaNet MSB-09151-01-11 0 0 0 0 0 0.250 MSB-09977-01-22 1 1 1 0 1 1 Her2Pos_Case_66 1 1 1 1 0.833 0.806 SISH + DenseNet MSB-09151-01-11 0 0 0 0 0 0 MSB-09977-01-22 0 0 0 0 0 0 Her2Pos_Case_66 0 0 0 0 0 0 RetCCL MSB-09151-01-11 0 0 0 1 0 0.250 MSB-09977-01-22 1 1 0 0 0 1 Her2Pos_Case_66 0 0 0 1 0 0.325 HSHR MSB-09151-01-11 1 1 1 1 0.833 0.833 MSB-09977-01-22 1 1 1 1 1 0.700 Her2Pos_Case_66 1 1 1 0 1 0.867
{stable}
[!ht] Evaluation results of different methods on Reader Study slides for subtype retrieval task. Method Reader Study MV@1 MV@3 MV@5 AP@3 AP@5 YOTTIXEL + KimiaNet MSB-09151-01-11 - - - - - MSB-09977-01-22 0 0 0 0 0 Her2Pos_Case_66 - - - - - SISH + DenseNet MSB-09151-01-11 - - - - - MSB-09977-01-22 0 0 0 0 0 Her2Pos_Case_66 - - - - - RetCCL MSB-09151-01-11 - - - - - MSB-09977-01-22 0 1 1 0.583 0.589 Her2Pos_Case_66 - - - - - HSHR MSB-09151-01-11 - - - - - MSB-09977-01-22 1 1 1 0.833 0.806 Her2Pos_Case_66 - - - - -
{stable}
[!h] Mann-Whitney U Test Results for Microscope Microscope Metric U-statistic P-value SISH MV_at_10_site nan nan AP_at_5_site 578.000 1.0000000000000000 MV_at_5_subtype nan nan AP_at_5_subtype 543.000 0.6156890923465373 HSHR MV_at_10_site 658.500 0.0670269674521155 AP_at_5_site 523.500 0.6135535212442216 MV_at_5_subtype 672.500 0.0905435676918205 AP_at_5_subtype 636.000 0.3353506099579945 Yottixel MV_at_10_site 523.500 0.5298120557113710 AP_at_5_site 455.500 0.1566946582170631 MV_at_5_subtype 532.000 0.6415761864226559 AP_at_5_subtype 534.500 0.7372510705087723
{stable}
[!h] Mann-Whitney U For Her2+ Model Metric U-statistic P-value Yottixel MV_at_10_site 2906.500 0.0000006008842402 AP_at_5_site 3287.500 0.0008328526912513 SISH MV_at_10_site nan nan AP_at_5_site 3651.500 0.0130448102685896 RetCCL MV_at_10_site nan nan AP_at_5_site 2483.500 0.0000000486711422 HSHR MV_at_10_site 3407.500 0.0002646668324846 AP_at_5_site 3274.500 0.0008856874714843
{sfigure}
[!ht] [Uncaptioned image] Performance comparison of Yottixel and SISH algorithms using Kima Net and Densenet on the Yale Trastuzumab dataset across various metrics.
{sfigure}
[!ht] [Uncaptioned image] Metrics for HSHR Microscope comparison analysis and Her2 analysis.
{sfigure}
[!ht] [Uncaptioned image] Metrics for RetCCL Her2 analysis.
{sfigure}
[!ht] [Uncaptioned image] Metrics for YOTTIXEL Her2 analysis.
{sfigure}
[!ht] [Uncaptioned image] Metrics for SISH Her2 analysis.